# Ergodic behaviors in reversible 3-state cellular automata

### v1

@@ 1 Introduction
The first attempt in this direction has been a classification of 2-state (binary) RCA with ‘lightcone’ or ‘relativistic’ local rules by Takesue [13] and Bobenko et al. [14]. Remarkably, among $2562^{8}=$ 256 rules one can identify a few exactly solvable rules which describe generic physical phenomena, such as diffusion, convection, etc. Examples are the so-called rules 54, 150, and 201 [15] (labeling the automata according to the binary code of the rule [14]). However, it has soon become clear that the set of effective dynamics of RCA can be much richer than those described by binary local rules (see e.g., [16, 17]). In this paper, we fill this observational gap by systematically exploring a set of one-dimensional 3-state RCA with all possible local rules (physical interactions) respecting certain physical symmetries. Specifically, we will focus on the so-called brickwork RCA, sometimes referred to as block cellular automata, where at each instant of time, a reversible mapping is applied to a block (brick) of two neighboring degrees of freedom $U:\mathbb{Z}_{k}\times\mathbb{Z}_{k}\rightarrow\mathbb{Z}_{k}\times\mathbb{Z}_{k}$, while in the following time instant, the reference point is shifted by one site (so after two time steps the reference point is again back at the original). The number of all reversible block cellular automata is equal to the number of reversible local maps, i.e., $(k^{2})!$. In the binary case, $k=2$, it turns out that dynamics generated by all such rules is rather uninteresting, with the least trivial example being perhaps the so-called deterministic East model [18]. Block RCA rules with $k>2$ have also been studied in specific cases, for example, see Ref. [17], while the entire set of block RCA is already prohibitively large to allow for a systematic study even for $k=3$.
@@
In the case $k=3$, which we systematically study in this work, this amounts to a total of $8!=40320$ rules. Identifying the remaining two states as particles with charge ±$\pm 1$, we shall focus on particular sets of symmetric block RCA sharing time-reversal, spatial-reversal (reflection) and charge-conjugation symmetries, or various combinations thereof. This then amounts to studying only a few hundred distinct symmetric BCA which we classify into 4 groups (classes) based on their fundamental dynamical (ergodic) and complexity behavior. Due to a fundamental difficulty with a proper definition of Lyapunov exponents and dynamical entropies for systems with discrete degrees of freedom, we determine the ergodicity class by a combination of the following empirical indicators: (i) the volume scaling of (average/typical) return time (starting from a random initial condition), (ii) temporal decay of spatio-temporal correlation functions of local observables (or densities of conserved charges), and (iii) leading eigenmodes of a dynamical transfer matrix of local (translationally invarant) observables. The latter correspond to either local (or quasilocal) conserved quantities or Ruelle-Pollicott resonances (chaotic modes), in gapless or gapped cases, respectively. We stress that this may be the first clear observation of the existence of Ruelle-Pollicott resonances in discrete state space dynamical systems and can serve as a definition of deterministic chaos.
@@ 2 Summary of results
The goal of this work is to propose a classification of different types of ergodic behavior in classical, time-reversible three-state cellular automata in one spatial dimension. We introduce this family of models by considering a periodic lattice of even length $L$. The microscopic state of the model is given by a configuration $s=(s_{1},s_{2},s_{3},\ldots,s_{L})$, where each $s_{j}$ takes a value from the local configuration space $X=\{+,-,\varnothing\}$. The setup can be seen as a toy model for a system of charged particles and anti-particles (with opposite charges $+$ and --) and vacuum $\varnothing$. We depict the sites with $s_{i}=+$ in red and the sites with $s_{i}=-$ in blue color. Absence of a ‘particle’ with $s_{i}=\varnothing$ (vacuum) is represented in white.
@@
Figure 1: The configuration-labeling scheme used in this work. From left to right, the sequence of integers from 00 to $8$ designates neighboring sites with all possible configurations from (,)($\varnothing$,$\varnothing$)( ∅ , ∅ ) to (+,)(+,-)( + , - ). White, blue and red colors are used to represent $\varnothing$, -- and $+$, respectively.
A rule that defines the evolution of the automaton is then given by a set of $8$ numbers specifying what other configuration each 2-site configuration in Figure 1 maps to. We assume that (,)($\varnothing$,$\varnothing$)( ∅ , ∅ ) always remains (,)($\varnothing$,$\varnothing$)( ∅ , ∅ ) and that the mapping rules at even and odd time steps are the same. Then, we use a shorthand notation with the sequence of eight numbers $N_{1}N_{2}\ldots N_{8}$ that designates what each configuration from Figure 1 (with the exception of the trivial 0, which always maps to 0) maps to. That is, $1\to N_{1}$, $2\to N_{2}$, and so on, until $8\to N_{8}$. For example, the rule $21354678$ means that under every step of discrete time evolution, $1\to 2$, $2\to 1$, $3\to 3$, $4\to 5$, $5\to 4$, $6\to 6$, $7\to 7$ and $8\to 8$. We show this example in Figure 2. Formal details of the setup and more mathematically precise definitions are given in Section 3.
@@
We note that there is a small number of models (in total 35 in all considered symmetry classes) for which it was not possible to clearly establish either one or more of the empirical criteria in a finite amount of accessible computation time. It is also possible that some of these models are fundamentally unclassifiable according to our scheme, e.g., their mean return time may be an erratic (or number theoretic) function of the system size. See Appendix C for details.
@@ 3 Three-state vacuum preserving cellular automata: models and symmetries
$\mathcal{V}_{o}=U_{2,3}\leavevmode\nobreak\ U_{4,5}\leavevmode\nobreak\ ...\leavevmode\nobreak\ U_{L,1},\qquad\mathcal{V}_{e}=U_{1,2}\leavevmode\nobreak\ U_{3,4}\leavevmode\nobreak\ ...\leavevmode\nobreak\ U_{L-1,L},$ (2)
@@ 3.1 Symmetries
[…]
@@ 3.2 Additional dynamical constraints
    $\forall s_{1},s_{2},s_{3},s_{4}\in X:\leavevmode\nobreak\ \leavevmode\nobreak\ U(s_{1},s_{2})=(s_{3},s_{4})\quad\text{and}\quad U_{D}(s_{1},s_{3})=(s_{2},s_{4}).$ (9)
@@
    $\text{East}:\quad\forall s_{1},s_{2},s_{3},s_{4}\in X:\leavevmode\nobreak\ \leavevmode\nobreak\ U(s_{1},s_{2})=(s_{3}=s_{1},s_{4}).$ (10)
@@
    $\text{West}:\quad\forall s_{1},s_{2},s_{3},s_{4}\in X:\leavevmode\nobreak\ \leavevmode\nobreak\ U(s_{1},s_{2})=(s_{3},s_{4}=s_{2}).$ (11)
@@ 4.1 Return time
[…]
@@
$X^{L}=\bigcup_{\alpha=1}^{N_{O}}\mathcal{O}_{s_{\alpha}},\leavevmode\nobreak\ \text{where}\leavevmode\nobreak\ s_{\beta}\notin\mathcal{O}_{s_{\alpha}}\leavevmode\nobreak\ \text{for}\leavevmode\nobreak\ \alpha\neq\beta,$ (13)
@@
[…]
@@
[…]
@@
We calculate average return times as a function of the number of sites $L$ for the BCA models given by the local update rules discussed above, see Figure 5. For generic models with the absence of conserved quantities, the average return time grows exponenially in system size $T_{U}=e^{\kappa L}$. In the majority of cases, the exponent parameter is submaximal $\kappa<\ln{3}$ ). It follows from Eqs. (16) and (15) that in such cases,
@@ 4.2 Conserved charges
[…]
@@
[…]
Note that all local generators add up to a unit element of the algebra (local decomposition of identity): $[+]_{x}+[-]_{x}+[\varnothing]_{x}\equiv\mathds{1}_{x}\equiv\mathds{1}$. Using the multiplicative property of the generators we then define the following generators
[…]
which spans the algebra $\mathcal{A}^{(r)}_{x}$ of all observables supported on the interval $[x,x+r-1]$ of length $r$:
[…]
@@
[…]
Notice that, due to brickwork geometry of our cellular automata setup, the map $\mathbb{U}$ is invariant up to translation with respect to two sites [^2] In fact, in the special case where $\mathcal{V}_{o}$ commutes with $\mathcal{V}_{e}$, the map $\mathcal{V}$ is invariant up to a translation of one site., i.e., $[\mathbb{U},\mathbb{T}^{2}]=0$. Thus we may consider dynamics confined to extensive observables invariant under a two-site translation:
@@
$\mathcal{A}_{\mathbb{T}}^{(r)}=\left\{\sum\limits_{j=1}^{L/2-1}\mathbb{T}^{2j}a\,|\,\leavevmode\nobreak\ a\in\mathcal{A}_{1}^{(r)}\right\}.$ (25)
@@
$A=\sum_{r}A^{(r)},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ A^{(r)}=P_{\mathcal{A}^{(r)}_{\mathbb{T}}/\mathcal{A}^{(r-1)}_{\mathbb{T}}}A,$ (26)
@@
$\mathbb{U}Q=Q\quad\Leftrightarrow\quad Q(\mathcal{V}s)=Q(s),\leavevmode\nobreak\ \forall s\in X^{L}.$ (27)
@@
    $Q^{(r)}=\sum_{j}\mathbb{T}^{2j}q^{(r)},\leavevmode\nobreak\ q^{(r)}\in\mathcal{A}^{(r)}_{1}.$ (28)
@@
    $0<\lim\limits_{L\rightarrow\infty}\frac{1}{L}\left<Q_{L},Q_{L}\right><\infty,$ (31)
    and has a fixed ($L$-independent) nonzero overlap with any strictly local observable (to set the scale). Notice that all quasilocal charges are pseudolocal, while, if $\|q^{(r)}\|\sim 1/r^{\alpha},\text{where}\leavevmode\nobreak\ \alpha>1$, the charge will be pseudolocal, but not quasilocal.
@@
    $\mathbb{U}^{n}Q=Q\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \mathbb{U}^{m}Q\neq Q,\leavevmode\nobreak\ \text{for}\leavevmode\nobreak\ 1\leq m<n.$ (32)
@@ 4.3 Transfer matrix of dynamics of observables
[…]
@@
Below, we will study the spectra of the transfer matrices $\mathcal{T}^{(r)}$ for increasing range $r$. Because we introduce the projection operator in the definition, the transfer matrix is not unitary, but is bounded in the operator norm $\|\mathcal{T}^{(r)}\|\equiv\sqrt{{\rm sup}_{A}\langle\mathcal{T}^{(r)}A,\mathcal{T}^{(r)}A\rangle/\langle A,A\rangle}\leq 1$. The eigenvalues of the transfer matrix $\mathcal{T}^{(r)}\mathcal{O_{\lambda}}=\lambda\leavevmode\nobreak\ \mathcal{O_{\lambda}}$ therefore lie on the unit disk $|\lambda|\leq 1$.
Figure 6: A sketch of the action of the transfer matrix $\mathcal{T}^{(r)}$ on the space of extensive local operators. We begin with an operator with support $r=3$, which is spread to support $5$ during a one-step time evolution $\mathbb{U}$, and is projected back to the $3$-local space as denoted by the red dashed lines.
@@
$\mathbb{U}Q^{(r^{\prime})}=Q^{(r^{\prime})}\leavevmode\nobreak\ \Rightarrow\leavevmode\nobreak\ \mathcal{T}^{(r)}Q^{(r^{\prime})}=Q^{(r^{\prime})},\leavevmode\nobreak\ \text{for}\leavevmode\nobreak\ r^{\prime}\leq r.$ (34)
@@
$\mathbb{U}^{n}Q^{(r^{\prime})}=Q^{(r^{\prime})}\leavevmode\nobreak\ \Rightarrow\leavevmode\nobreak\ \left[\mathcal{T}^{(r)}\right]^{n}Q^{(r)}=Q^{(r)},\leavevmode\nobreak\ \textrm{for}\leavevmode\nobreak\ r\geq r^{\prime}+2(n-1).$ (35)

### v2

@@ 1 Introduction
The first attempt in this direction has been a classification of 2-state (binary) RCA with ‘lightcone’ or ‘relativistic’ local rules by Takesue [13] and Bobenko et al. [14]. Remarkably, among $2^{8}=256$ rules one can identify a few exactly solvable rules which describe generic physical phenomena, such as diffusion, convection, etc. Examples are the so-called rules 54, 150, and 201 [15] (labeling the automata according to the binary code of the rule [14]). However, it has soon become clear that the set of effective dynamics of RCA can be much richer than those described by binary local rules (see e.g., [16, 17]). In this paper, we fill this observational gap by systematically exploring a set of one-dimensional 3-state RCA with all possible local rules (physical interactions) respecting certain physical symmetries. Specifically, we will focus on the so-called brickwork RCA, sometimes referred to as block cellular automata, where at each instant of time, a reversible mapping is applied to a block (brick) of two neighboring degrees of freedom $U:\mathbb{Z}_{k}\times\mathbb{Z}_{k}\rightarrow\mathbb{Z}_{k}\times\mathbb{Z}_{k}$, while in the following time instant, the reference point is shifted by one site (so after two time steps the reference point is again back at the original). The number of all reversible block cellular automata is equal to the number of reversible local maps, i.e., $(k^{2})!$. In the binary case, $k=2$, it turns out that dynamics generated by all such rules is rather uninteresting, with the least trivial example being perhaps the so-called deterministic East model [18]. Block RCA rules with $k>2$ have also been studied in specific cases, for example, see Ref. [17], while the entire set of block RCA is already prohibitively large to allow for a systematic study even for $k=3$.
@@
In the case $k=3$, which we systematically study in this work, this amounts to a total of $8!=40320$ rules. Identifying the remaining two states as particles with charge $\pm 1$, we shall focus on particular sets of symmetric block RCA sharing time-reversal, spatial-reversal (reflection) and charge-conjugation symmetries, or various combinations thereof. This then amounts to studying only a few hundred distinct symmetric BCA which we classify into 4 groups (classes) based on their fundamental dynamical (ergodic) and complexity behavior. Due to a fundamental difficulty with a proper definition of Lyapunov exponents and dynamical entropies for systems with discrete degrees of freedom, we determine the ergodicity class by a combination of the following empirical indicators: (i) the volume scaling of (average/typical) return time (starting from a random initial condition), (ii) temporal decay of spatio-temporal correlation functions of local observables (or densities of conserved charges), and (iii) leading eigenmodes of a dynamical transfer matrix of local (translationally invarant) observables. The latter correspond to either local (or quasilocal) conserved quantities or Ruelle-Pollicott resonances (chaotic modes), in gapless or gapped cases, respectively. We stress that this may be the first clear observation of the existence of Ruelle-Pollicott resonances in discrete state space dynamical systems and can serve as a definition of deterministic chaos.
@@ 2 Summary of results
The goal of this work is to propose a classification of different types of ergodic behavior in classical, time-reversible three-state cellular automata in one spatial dimension. We introduce this family of models by considering a periodic lattice of even length $L$. The microscopic state of the model is given by a configuration $s=(s_{1},s_{2},s_{3},\ldots,s_{L})$, where each $s_{j}$ takes a value from the local configuration space $X=\{+,-,\varnothing\}$. The setup can be seen as a toy model for a system of charged particles and anti-particles (with opposite charges ++ and -) and vacuum $\varnothing$. We depict the sites with $s_{i}=+$ in red and the sites with $s_{i}=$- in blue color. Absence of a ‘particle’ with $s_{i}=\varnothing$ (vacuum) is represented in white.
@@
Figure 1: The configuration-labeling scheme used in this work. From left to right, the sequence of integers from 0 to 88 designates neighboring sites with all possible configurations from $(\varnothing,\varnothing)$ to (+,)(+,-). White, blue and red colors are used to represent $\varnothing$, - and ++, respectively.
A rule that defines the evolution of the automaton is then given by a set of $8$ numbers specifying what other configuration each 2-site configuration in Figure 1 maps to. We assume that $(\varnothing,\varnothing)$ always remains $(\varnothing,\varnothing)$ and that the mapping rules at even and odd time steps are the same. Then, we use a shorthand notation with the sequence of eight numbers $N_{1}N_{2}\ldots N_{8}$ that designates what each configuration from Figure 1 (with the exception of the trivial 0, which always maps to 0) maps to. That is, $1\to N_{1}$, $2\to N_{2}$, and so on, until $8\to N_{8}$. For example, the rule $21354678$ means that under every step of discrete time evolution, $1\to 2$, $2\to 1$, $3\to 3$, $4\to 5$, $5\to 4$, $6\to 6$, $7\to 7$ and $8\to 8$. We show this example in Figure 2. Formal details of the setup and more mathematically precise definitions are given in Section 3.
@@
We note that there is a small number of models (in total 35 in all considered symmetry classes) for which it was not possible to clearly establish either one or more of the empirical criteria in a finite amount of accessible computation time. It is also possible that some of these models are fundamentally unclassifiable according to our scheme, e.g., their mean return time may be an erratic (or number theoretic) function of the system size. See Appendix B for details.
@@ 3 Three-state vacuum preserving cellular automata: models and symmetries
$\mathcal{V}_{o}=U_{2,3}\penalty 10000\ U_{4,5}\penalty 10000\ ...\penalty 10000\ U_{L,1},\qquad\mathcal{V}_{e}=U_{1,2}\penalty 10000\ U_{3,4}\penalty 10000\ ...\penalty 10000\ U_{L-1,L},$ (2)
@@ 3.1 Symmetries
    $\mathcal{C}(s_{1},s_{2})\overset{\text{def}}{=}(s_{1}^{*},s_{2}^{*}),\penalty 10000\ \text{where}\penalty 10000\ s^{*}=\begin{cases}-,\penalty 10000\ \penalty 10000\ \text{if}\penalty 10000\ s=+\\ +,\penalty 10000\ \penalty 10000\ \text{if}\penalty 10000\ s=-\\ \varnothing,\penalty 10000\ \penalty 10000\ \text{if}\penalty 10000\ s=\varnothing\end{cases}.$ (6)
@@ 3.2 Additional dynamical constraints
    $\forall s_{1},s_{2},s_{3},s_{4}\in X:\penalty 10000\ \penalty 10000\ U(s_{1},s_{2})=(s_{3},s_{4})\quad\text{and}\quad U_{D}(s_{1},s_{3})=(s_{2},s_{4}).$ (9)
@@
    $\text{East}:\quad\forall s_{1},s_{2},s_{3},s_{4}\in X:\penalty 10000\ \penalty 10000\ U(s_{1},s_{2})=(s_{3}=s_{1},s_{4}).$ (10)
@@
    $\text{West}:\quad\forall s_{1},s_{2},s_{3},s_{4}\in X:\penalty 10000\ \penalty 10000\ U(s_{1},s_{2})=(s_{3},s_{4}=s_{2}).$ (11)
@@ 4.1 Return time
$\mathcal{O}_{s}\overset{\text{def}}{=}\{s^{\prime}\in X^{L}\,|\,\exists t:\penalty 10000\ s^{\prime}=\mathcal{V}^{t}s\},$ (12)
@@
$X^{L}=\bigcup_{\alpha=1}^{N_{O}}\mathcal{O}_{s_{\alpha}},\penalty 10000\ \text{where}\penalty 10000\ s_{\beta}\notin\mathcal{O}_{s_{\alpha}}\penalty 10000\ \text{for}\penalty 10000\ \alpha\neq\beta,$ (13)
@@
$T_{U}\overset{\text{def}}{=}\frac{1}{3^{L}}\sum\limits_{s\in X^{L}}|\mathcal{O}_{s}|=\frac{1}{3^{L}}\sum\limits_{\alpha=1}^{N_{O}}|\mathcal{O}_{s_{\alpha}}|^{2},$ (14)
@@
$\frac{{\mathcal{O}}_{max}}{3^{L}}\leq\left(\frac{T_{U}}{3^{L}}\right)^{1/2},\penalty 10000\ \text{where}\penalty 10000\ \penalty 10000\ {\mathcal{O}}_{max}\overset{\text{def}}{=}\max\limits_{\alpha}|\mathcal{O}_{s_{\alpha}}|.$ (15)
@@
[…]
@@ 4.2 Conserved charges
Observables. A global observable $\mathcal{O}:X^{L}\rightarrow\mathbb{C}$ is a complex-valued function in the space of configurations. The global observables form a multiplicative algebra $\mathcal{A}$ of functions over the configuration space. Multiplication is given by $\left(\mathcal{O}_{1}\bm{\cdot}\mathcal{O}_{2}\right)(s)=\mathcal{O}_{1}(s)\mathcal{O}_{2}(s)$, so the algebra is commutative. The algebra $\mathcal{A}$ is also a vector space, allowing us to define the inner product $\left<\penalty 10000\ ,\penalty 10000\ \right>:\mathcal{A}\times\mathcal{A}\rightarrow\mathbb{C}$:
$\left<\mathcal{O}_{1},\mathcal{O}_{2}\right>\overset{\text{def}}{=}\frac{1}{|X^{L}|}\sum_{s}\mathcal{O}^{*}_{1}(s)\mathcal{O}_{2}(s).$ (18)
@@
$\alpha]_{x}(s)\overset{\text{def}}{=}\delta_{\alpha,s_{x}},\penalty 10000\ \penalty 10000\ \penalty 10000\ \alpha\in X.$ (19)
Note that all local generators add up to a unit element of the algebra (local decomposition of identity): $_{x}+[-]_{x}+[\varnothing]_{x}\equiv\mathds{1}_{x}\equiv\mathds{1}$. Using the multiplicative property of the generators we then define the following generators
$\alpha_{1}\alpha_{2}\ldots\alpha_{r}]_{x}\overset{\text{def}}{=}[\alpha_{1}]_{x}\bm{\cdot}[\alpha_{2}]_{x+1}...\bm{\cdot}[\alpha_{r}]_{x+r-1},$ (20)
which spans the algebra $\mathcal{A}^{(r)}_{x}$ of all observables supported on the interval [x,x+r1][x,x+r-1] of length $r$:
$\mathcal{A}_{x}^{(r)}\overset{\text{def}}{=}\text{span}\{[\alpha_{1}\alpha_{2}\ldots\alpha_{r}]_{x}\,|\,\alpha_{i}\in X\},\penalty 10000\ x=1,2,\ldots,L.$ (21)
@@
$\mathbb{U}\mathcal{O}(s)\overset{\text{def}}{=}\mathcal{O}\left(\mathcal{V}s\right),\quad\mathcal{O}^{t}\overset{\text{def}}{=}\mathbb{U}^{t}\mathcal{O}.$ (23)
Notice that, due to brickwork geometry of our cellular automata setup, the map $\mathbb{U}$ is invariant up to translation with respect to two sites [^2] In fact, in the special case where $\mathcal{V}_{o}$ commutes with $\mathcal{V}_{e}$, the map $\mathcal{V}$ is invariant up to a translation of one site., i.e., $\mathbb{U},\mathbb{T}^{2}]=0$. Thus we may consider dynamics confined to extensive observables invariant under a two-site translation:
@@
$\mathcal{A}_{\mathbb{T}}^{(r)}=\left\{\sum\limits_{j=1}^{L/2-1}\mathbb{T}^{2j}a\,|\,\penalty 10000\ a\in\mathcal{A}_{1}^{(r)}\right\}.$ (25)
@@
$A=\sum_{r}A^{(r)},\penalty 10000\ \penalty 10000\ \penalty 10000\ A^{(r)}=P_{\mathcal{A}^{(r)}_{\mathbb{T}}/\mathcal{A}^{(r-1)}_{\mathbb{T}}}A,$ (26)
@@
$\mathbb{U}Q=Q\quad\Leftrightarrow\quad Q(\mathcal{V}s)=Q(s),\penalty 10000\ \forall s\in X^{L}.$ (27)
@@
    $Q^{(r)}=\sum_{j}\mathbb{T}^{2j}q^{(r)},\penalty 10000\ q^{(r)}\in\mathcal{A}^{(r)}_{1}$. (28)
@@
    $0<\lim\limits_{L\rightarrow\infty}\frac{1}{L}\left<Q_{L},Q_{L}\right$$\infty$, (31)
    and has a fixed ($L$-independent) nonzero overlap with any strictly local observable (to set the scale). Notice that all quasilocal charges are pseudolocal, while, if $\|q^{(r)}\|\sim 1/r^{\alpha},\text{where}\penalty 10000\ \alpha>1$, the charge will be pseudolocal, but not quasilocal.
@@
    $\mathbb{U}^{n}Q=Q\penalty 10000\ \text{and}\penalty 10000\ \mathbb{U}^{m}Q\neq Q,\penalty 10000\ \text{for}\penalty 10000\ 1\leq m<n.$ (32)
@@ 4.3 Transfer matrix of dynamics of observables
$\mathcal{T}^{(r)}\overset{\text{def}}{=}P_{\mathcal{A}_{\mathbb{T}}^{(r)}}\,\cdot\,\mathbb{U},$ (33)
@@
Below, we will study the spectra of the transfer matrices $\mathcal{T}^{(r)}$ for increasing range $r$. Because we introduce the projection operator in the definition, the transfer matrix is not unitary, but is bounded in the operator norm $\|\mathcal{T}^{(r)}\|\equiv\sqrt{{\rm sup}_{A}\langle\mathcal{T}^{(r)}A,\mathcal{T}^{(r)}A\rangle/\langle A,A\rangle}\leq 1$. The eigenvalues of the transfer matrix $\mathcal{T}^{(r)}\mathcal{O_{\lambda}}=\lambda\penalty 10000\ \mathcal{O_{\lambda}}$ therefore lie on the unit disk $|\lambda|\leq 1$.
Figure 6: A sketch of the action of the transfer matrix $\mathcal{T}^{(r)}$ on the space of extensive local operators. We begin with an operator with support $r=3$, which is spread to support 55 during a one-step time evolution $\mathbb{U}$, and is projected back to the 33-local space as denoted by the red dashed lines.
@@
$\mathbb{U}Q^{(r^{\prime})}=Q^{(r^{\prime})}\penalty 10000\ \Rightarrow\penalty 10000\ \mathcal{T}^{(r)}Q^{(r^{\prime})}=Q^{(r^{\prime})},\penalty 10000\ \text{for}\penalty 10000\ r^{\prime}\leq r.$ (34)
@@
$\mathbb{U}^{n}Q^{(r^{\prime})}=Q^{(r^{\prime})}\penalty 10000\ \Rightarrow\penalty 10000\ \left[\mathcal{T}^{(r)}\right]^{n}Q^{(r)}=Q^{(r)},\penalty 10000\ \textrm{for}\penalty 10000\ r\geq r^{\prime}+2(n-1).$ (35)